OpenStudy (anonymous):

The perimeter of a triangle is 24x+35 . One side is 8x-9 and the other side is 12x+20 . Find the missing side.

ganeshie8 (ganeshie8):

perimeter = side1 + side2 + side3 24x + 35 = 8x-9 + 12x+20 + side3 24x + 35 = 20x + 11 + side3 4x+24 = side3

OpenStudy (anonymous):

thanks hun!

ganeshie8 (ganeshie8):

np yw :)